# Tactile-based Object Retrieval From Granular Media

### v2

@@ III-C Reward Design and Policy Architecture
Our policy has one actor network and one critic network, trained with PPO [70]. Both networks are a 5-layer multilayer perceptron (MLP), and each layer has 256 hidden neurons. The weights are not shared between the actor and critic. We use arctan as our activation function between layers. We note that we also tried a Transformer [71] and LSTM [72] encoder to learn the relationship among observations from different steps in our history buffer; however, we found the fully connected MLP learns faster and converges to a higher reward.
III-D Curriculum Training Strategy
Figure 3: Curriculum training strategy. We first pre-train our policy on a tabletop environment and then fine-tune it inside granular media. We then zero-shot deploy our policy on the real hardware for evaluation. This curriculum allows the policy to converge to a higher performance within a shorter training time.
@@ III-D Curriculum Training Strategy
IV Experimental Setup
IV-A Real-world Setup
We use a WSG50 parallel gripper with two DISCO [7] tactile fingers, as shown in Fig. 4. We mount this gripper to a UR5 robotic arm to control its movement. This setup is effective for interacting with objects under granular materials, as the hemispherical top allows easy penetration into granular media and the cylindrical shape allows smooth movement with small drag. This multi-curved tactile finger has sensing abilities covering the hemisphere top and cylinder. This all-around sensing coverage allows it to sense contact from many directions, which is essential for exploration under granular media. When fully opened, the distance between the two finger centers is 143mm.
@@ IV-A Real-world Setup
Figure 5: 35 objects used in our real-robot experiments. Our method is trained on a set of seven objects. It is then evaluated quantitatively (ten attempts on each object) on both the set of seven training objects and a set of six previously unseen objects. In additional testing, we also show that our method can work on an extended set of 22 unseen objects with at most three attempts per object.
@@
The DISCO finger returns a single contact containing the contact location vector and force vector. To filter out ubiquitous contacts with the granular media and keep only the pushing contacts with the target object, we discard those contacts with a force magnitude smaller than 4N. We fill in zeros into the location and force vectors to our policy when there are no pushing contacts detected at a particular step. We find the 4N filtering threshold empirically through calibration, and we use the same 4N threshold for all objects. The optimal filtering threshold should vary across different objects and different types of granular media. But our policy learns to be robust to spurious pushing contacts so we do not need object and granular media-specific calibration. The granular media in our experiments are pony beads with a diameter of 9mm. Our gripper with DISCO fingers is able to move smoothly within such media without jamming.
@@
In our real-world experiments, we evaluate our method on 35 objects in total, as shown in Fig. 5. Our object set covers a variety of complex shapes, including rigid, deformable, and articulated objects. The object weight in our set also varies a lot, with the lightest object (pen) being 4.5g and the heaviest object (sugar box) being 470g. We evaluate quantitatively on 13 objects with 10 trials each and among them, 7 objects are used for training in simulation and 6 objects are unseen. We then qualitatively show that our method can work on another 22 unseen objects randomly selected from our lab, each within at most three attempts.
@@
The gripper moves straight forward in the beginning until the first pushing contact, and then our learned policy starts guiding the exploration. We bury the object such that the center of the object is roughly in front of one of the fingers. This is to guarantee an initial first contact as the finger moves forward. Between each trial, we alternate the location of the object in front of the left and right fingers, and we also rotate the object to cover a diverse set of orientations. The episode starts with the first contact and then ends when the maximum steps of 100 are reached or the policy executes a grasp.
IV-B Simulation Setup
@@
Different from the real-world experiments where we only evaluate inside granular media, we have two different setups for the curriculum training strategy in simulation: tabletop and granular media. We build the two environments using the IsaacGym physics simulator [73]. For both environments, we simulate a floating parallel gripper that moves in a 2D plane, according to its real hardware dimensions.
@@
Despite the fact that we attempt to bury the object such that its center is in front of one of the fingers, it is very hard to guarantee it in practice on the real setup. As a result, in order to be robust to the object location variance, we add more randomization on the object location and orientation during training in simulation. In each episode in the simulation, the object is loaded at a fixed distance away from the gripper. The exact distance does not matter since the finger always moves straight forward until the first contact. However, we vary the location of the object along the direction perpendicular to the gripper’s approaching direction by ±$\pm 2$cm. We also randomize the orientation of the object. It is possible that for small objects like pentagrams, the finger can pass the object without making contact under the location randomization, in which case we simply move on to the next episode.
@@
IV-C Sensor Noise
Real-world tactile sensors all show some level of erroneous readings, especially inside granular media. One advantage of the learning-based method is that we can incorporate noise into our training process in simulation for more robust transfer onto the real robot. For a single contact on our DISCO finger, the reported level for the contact location noise is between 1mm and 2mm on average, and the reported level for the contact force noise is 0.1N on average for a force level of $3\sim 4$N. As a result, due to increased noise level in granular media, we apply a larger location noise uniformly distributed from $[-1\text{cm},1\text{cm}]$ and a larger force noise uniformly distributed from $[-0.2\text{N},0.2\text{N}]$ to each element in the contact location and net force vectors during training in simulation for all learning-based baselines.
IV-D Experimental Conditions and Baselines
@@
IV-D1 Align-to-grasp (A2G)
This method rotates the gripper to align with the direction of the contact force after the first pushing contact is detected. This is a heuristics-based method and does not require training, intended to evaluate the importance of a learning-based method that trains in the presence of noise.
IV-D2 A2G-push
This variant of A2G starts rotating and aligning only after five consecutive contacts on the same finger have been detected. This is intended to filter out spurious pushing contacts in granular media, and to evaluate the performance of pushing alone as a heuristic for reducing uncertainty.
IV-D3 GEOTACT-tabletop
This is the policy from the first stage of our curriculum training pipeline, trained on the open-air tabletop environment.
IV-D4 GEOTACT-granular
This is the policy defined in our method, but trained exclusively in granular media, thus bypassing our proposed curriculum.
IV-D5 GEOTACT
This is our full proposed method, combining all the aspects discussed in this study.
V Results and Discussion
@@
V-A Real-robot Results
TABLE I: Real-world experiment results. The number of successes out of 10 trials is shown for 7 training objects and 6 unseen objects.
Method Training Objects Unseen Objects
Cube Long Cuboid Ball Pentagram L-shape Tomato Soup Can Potted Meat Can Clamp Toy Banana Utility Knife Screw Driver Chocolate Box Coconut Water Can
A2G 3/10 3/10 0/10 2/10 4/10 1/10 4/10 3/10 3/10 3/10 2/10 4/10 4/10
GEOTACT 7/10 8/10 1/10 6/10 6/10 3/10 7/10 8/10 6/10 6/10 5/10 7/10 6/10
We evaluate two representative methods on the real robot - our method, GEOTACT, which is learning-based and uses pushing, and A2G, which is heuristics-based and does not use pushing. We first evaluate both methods on the training set (seven objects) and unseen set (six objects) shown in Fig. 5, performing ten trials on each object. The results are shown in Table I.
GEOTACT achieves a success rate of over 60% on a majority of objects and generalizes well to unseen objects without a noticeable performance drop. It outperforms A2G, which suffers from the significant amount of sensor noise introduced by granular media. In comparison, GEOTACT is more robust to sensing noise by continuously interacting with the object through emergent pushing actions and reacting to new contact observations. In addition, fine-tuned in granular media, GEOTACT is more adapted to the object motion discrepancy between tabletop and granular media. We also notice that the success rates of GEOTACT on two predominantly rounded shapes (ball and tomato soup can) are significantly lower compared to other objects. We discuss these failure cases, along with failure cases of A2G, later in Sec. V-C.
@@ V-A Real-robot Results
V-B Simulation Results
TABLE II: Simulation experiment results on the tabletop. Each success rate is computed over 1000 trials.
@@ V-B Simulation Results
TABLE III: Simulation experiment results inside granular materials. Each success rate is computed over 1000 trials.
@@
Figure 10: Bird’s-eye view of four failure scenarios for A2G. A2G consists of three stages: obtaining the first contact, aligning the gripper with the contact normal, and then closing the gripper. See Sec. V-C for a detailed explanation of these failure cases.
First, in an open-air tabletop environment, we evaluate A2G and GEOTACT-tabletop using the same amount of noise as used in training (shown in Table II). (Note that we do not evaluate A2G-push on the tabletop, as A2G-push is designed to be robust to spurious contact readings, which do not happen in an open-air environment.) Without the extra sensor noise and complex object motion introduced by granular media, GEOTACT-tabletop is achieving over 90% success rate. A2G also has a higher success rate (65%) compared to its evaluation in granular media on the real robot. However, A2G struggles with a few objects, such as L-shape, long cuboid, and pentagram, even in the tabletop environment, and we discuss these failure cases in detail in the next section.
Inside granular media simulation, we evaluate all five methods using the same amount of tactile noise as training (Table III). Working within granular media is much more challenging than the tabletop setup. As a result, the performance of both GEOTACT-tabletop and A2G drops significantly compared to tabletop simulation, from 91% to 27% and from 65% to 39%, respectively. A2G-push performs slightly better (44%) than A2G, showing that adding heuristic push can reduce uncertainty and increase robustness.
@@
V-C Failure Cases and Discussion
@@
VI Conclusion
@@
In this work, we propose GEOTACT, a novel method that reliably retrieves objects buried completely under granular media using only touch sensing. This is a challenging task due to the absolute absence of vision and spurious and noisy tactile readings, introduced by the ubiquitous contacts inside granular media. We formulate this task as a model-free reinforcement learning problem and train it end-to-end in simulation with a curriculum strategy. We first train the policy in an open-air environment and then fine-tune it inside a granular environment. Our formulation of the learning task action space leads to emergent pushing behaviors that help reduce uncertainty and also move the object to form a stable grasp. Compared to heuristics-based methods, our learned policy can incorporate simulated sensor noise during training and is thus more robust to spurious and noisy tactile readings. Despite being trained only on 7 objects with simple shapes in simulation, we show that our method can zero-shot transfer to the real hardware and successfully pick up 28 unseen objects, including rigid, deformable, and articulated objects with various complex shapes.
Acknowledgments
@@
We would like to thank Eric Chang and Amr El-Azizi for their support on the DISCO tactile finger. We would like to thank Zhenjia Xu for his help with the granular media simulation. This material is based upon work supported by the National Science Foundation under Award No. CMMI-2037101, and the Office of Naval Research under grant N00014-21-1-4010. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation or the Office of Naval Research.

### v3

@@ 3.3 Reward Design and Policy Architecture
Our policy has one actor network and one critic network, trained with PPO [schulman2017proximal]. Both networks are a 5-layer multilayer perceptron (MLP), and each hidden layer has 256 neurons. Since our task is not image-based, the actor and critic networks do not share any weights or layers, which is a common practice and shown to have better performance for control tasks [andrychowicz2020matters]. We use arctan as our activation function between layers. Notice that we also tried a Transformer [vaswani2017attention] and an LSTM [hochreiter1997long] encoder to learn the relationship among observations from different steps in our history buffer; however, we found that the fully connected MLP learns faster and converges to a higher reward.
3.4 Curriculum Training Strategy
@@
4 Experimental Setup
4.1 Real-world Setup
We use a WSG50 parallel gripper with two DISCO [piacenza2020sensorized] tactile fingers, as shown in Fig. 4. We mount this gripper to a UR5 robotic arm to control its movement. This setup is effective for interacting with objects under granular materials, as the hemispherical top allows easy penetration into granular media and the cylindrical shape allows smooth movement with small drag. This multi-curved tactile finger has sensing abilities covering the hemisphere top and cylinder. This all-around sensing coverage allows it to sense contact from many directions, which is essential for exploration under granular media. When fully opened, the distance between the two finger centers is 143mm.
@@ 4.1 Real-world Setup
Figure 5: 35 objects used in our real-robot experiments. Our method is trained on a set of 7 objects. It is then evaluated quantitatively (10 attempts on each object) on both the set of 7 training objects and a set of 6 previously unseen objects. In additional testing, we also show that our method can work on an extended set of 22 unseen objects with at most three attempts per object. The images of training and unseen objects do not correctly reflect their relative sizes. See the demonstration videos for each object on our website for accurate size information.
The DISCO finger returns a single contact containing the contact location vector and force vector. The tactile finger uses a pre-trained machine learning model to predict a single contact from the high-dimensional raw signals. When there are multiple contacts, the output contact location roughly matches the location of the contact with the largest force, and the output force roughly matches the combined net force from all contact forces.
@@
To filter out ubiquitous contacts with the granular media and keep only the pushing contacts with the target object, we discard those contacts with a force magnitude smaller than 4N. We fill in zeros into the location and force vectors to our policy when there are no pushing contacts detected at a particular step. We find the 4N filtering threshold empirically through calibration, and we use the same 4N threshold for all objects. The optimal filtering threshold should vary across different objects and different types of granular media. But our policy learns to be robust to spurious pushing contacts so we do not need object and granular media-specific calibration. The granular media in our experiments are pony beads with a diameter of 9mm. Our gripper with DISCO fingers is able to move smoothly within such media without jamming.
@@
In our real-world experiments, we evaluate our method on 35 objects in total, as shown in Fig. 5. Our object set covers a variety of complex shapes, including rigid, deformable, and articulated objects. The object weight in our set also varies a lot, with the lightest object (pen) being 4.5g and the heaviest object (sugar box) being 470g. We evaluate quantitatively on 13 objects with 10 trials each and among them, 7 objects are used for training in simulation and 6 objects are unseen. We then qualitatively show that our method can work on another 22 unseen objects randomly selected from our lab. We quickly go through this set of objects, and we move on to the next object as soon as a successful trial is achieved on the current object.
The gripper starts initially in the air, and it follows a predefined heuristic trajectory to move into the granular media, makes a shaking motion to shake off jammed beads, and then moves straight forward until the first pushing contact. We start applying force filtering after the finger starts moving forward inside the granular media to filter out ubiquitous contacts, and the learned policy only takes over after the first pushing contact is obtained.
@@
We bury the object so that part of the object occupies the initial forward trajectory of the finger. This is to guarantee an initial first contact as the finger moves forward. Between each trial, we randomize the location of the object, alternate the location of the object in front of the left and right fingers, and we also rotate the object to cover a diverse set of orientations. The episode starts with the first pushing contact and then ends when the maximum steps of 100 are reached or the policy executes a grasp.
4.2 Simulation Setup
@@
Different from the real-world experiments where we only evaluate inside granular media, we have two different setups for the curriculum training strategy in simulation: tabletop and granular media. We build the two environments using the IsaacGym physics simulator [makoviychuk2021isaac]. For both environments, we simulate a floating parallel gripper that moves in a 2D plane, according to its real hardware dimensions.
@@
Despite the fact that we attempt to bury the object such that its center is in front of one of the fingers, it is very hard to guarantee it in practice on the real setup. As a result, in order to be robust to the object location variance, we add more randomization on the object location and orientation during training in simulation. In each episode in the simulation, the object is loaded at a fixed distance away from the gripper. The exact distance does not matter since the finger always moves straight forward until the first contact. However, we vary the location of the object along the direction perpendicular to the gripper’s approaching direction by $\pm 2$cm. We also randomize the orientation of the object. It is possible that for small objects like pentagrams, the finger can pass the object without making contact under the location randomization, in which case we simply move on to the next episode.
@@
4.3 Sensor Noise
Real-world tactile sensors all show some level of erroneous readings, especially inside granular media. One advantage of the learning-based method is that we can incorporate noise into our training process in simulation for more robust transfer onto the real robot. For a single contact on our DISCO finger, the reported level for the contact location noise is between 1mm and 2mm on average, and the reported level for the contact force noise is 0.1N on average for a force level of $3\sim 4$N. As a result, due to increased noise level in granular media, we apply a larger location noise uniformly distributed from $1\text{cm},1\text{cm}]$ and a larger force noise uniformly distributed from $0.2\text{N},0.2\text{N}]$ to each element in the contact location and net force vectors during training in simulation for all learning-based baselines.
4.4 Experimental Conditions and Baselines
@@
Align-to-grasp (A2G). This method rotates the gripper to align with the direction of the contact force after the first pushing contact is detected. This is a heuristics-based method and does not require training, intended to evaluate the importance of a learning-based method that trains in the presence of noise.
A2G-push. This variant of A2G starts rotating and aligning only after five consecutive contacts on the same finger have been detected. This is intended to filter out spurious pushing contacts in granular media, and to evaluate the performance of pushing alone as a heuristic for reducing uncertainty.
GEOTACT-tabletop. This is the policy from the first stage of our curriculum training pipeline, trained on the open-air tabletop environment.
GEOTACT-granular. This is the policy defined in our method, but trained exclusively in granular media, thus bypassing our proposed curriculum.
GEOTACT. This is our full proposed method, combining all the aspects discussed in this study.
5 Results and Discussion
@@
5.1 Real-robot Results
Table 1: Real-world experiment results on 7 training objects. We show the number of successes out of 10 trials.
Method Training Objects Average
A2G 3/10 3/10 0/10 2/10 4/10 1/10 4/10 0.24
GEOTACT 7/10 8/10 1/10 6/10 6/10 3/10 7/10 0.54
Table 2: Real-world experiment results on 6 unseen objects. We show the number of successes out of 10 trials.
Method Unseen Objects Average
Clamp Toy Banana Utility Knife Screw Driver Chocolate Box Coconut Water Can
A2G 3/10 3/10 3/10 2/10 4/10 4/10 0.32
GEOTACT 8/10 6/10 6/10 5/10 7/10 6/10 0.63
We evaluate two representative methods on the real robot — our method, GEOTACT, which is learning-based and uses pushing, and A2G, which is heuristics-based and does not use pushing. We first evaluate both methods on the training set (seven objects) and unseen set (six objects) shown in Fig. 5, performing ten trials on each object. The results on the training set and the unseen set are shown in Table 1 and Table 2, respectively.
GEOTACT achieves success rates of 54% and 63% on the training set and unseen set, respectively, demonstrating strong generalization to unseen objects without any performance drop. It outperforms A2G, which suffers from the significant amount of sensor noise introduced by granular media. In comparison, GEOTACT is more robust to sensing noise by continuously interacting with the object through emergent pushing actions and reacting to new contact observations. In addition, fine-tuned in granular media, GEOTACT is more adapted to the object motion discrepancy between tabletop and granular media. We also notice that the success rates of GEOTACT on two predominantly rounded shapes (ball and tomato soup can) are significantly lower compared to other objects. We discuss these failure cases, along with failure cases of A2G, later in Sec. 5.3.
@@ 5.1 Real-robot Results
5.2 Simulation Results
Table 3: Simulation experiment results on the tabletop. Each success rate is computed over 1000 trials.
@@ 5.2 Simulation Results
Table 4: Simulation experiment results inside granular materials. Each success rate is computed over 1000 trials.
@@
First, in an open-air tabletop environment, we evaluate A2G and GEOTACT-tabletop using the same amount of noise as used in training (shown in Table 3). Note that we do not evaluate A2G-push on the tabletop, as A2G-push is designed to be robust to spurious contact readings, which do not happen in an open-air environment. Without the extra sensor noise and complex object motion introduced by granular media, GEOTACT-tabletop is achieving over 90% success rate. A2G also has a higher success rate (65%) compared to its evaluation in granular media on the real robot. However, A2G struggles with a few objects, such as L-shape, long cuboid, and pentagram, even in the tabletop environment, and we discuss these failure cases in detail in the next section.
Inside granular media simulation, we evaluate all five methods using the same amount of tactile noise as training (Table 4). Working within granular media is much more challenging than the tabletop setup. As a result, the performance of both GEOTACT-tabletop and A2G drops significantly compared to tabletop simulation, from 91% to 27% and from 65% to 39%, respectively. A2G-push performs slightly better (44%) than A2G, showing that adding heuristic push can reduce uncertainty and increase robustness.
@@
5.3 Failure Cases and Discussion
Figure 10: Bird’s-eye view of four failure scenarios for A2G. A2G consists of three stages: obtaining the first contact, aligning the gripper with the contact normal, and then closing the gripper. See Sec. 5.3 for a detailed explanation of these failure cases.
@@ 5.3 Failure Cases and Discussion
The two objects, ball and tomato soup can, is exactly the reason why our method has a higher success rate on unseen objects than training objects. Despite the fact that these two objects are seen during training, they are inherently hard to grasp due to the limitation of the parallel gripper with cylindrical fingers. If we exclude these two objects, the remaining objects in the training set actually have a higher average success rate than unseen objects. Most of our unseen objects have irregular shapes and are not cylindrical or spherical, which is not challenging from a hardware perspective. The coconut water can was originally cylindrical, but it was squeezed and lost its perfect cylindrical shape.
@@
6 Limitations and Future Works
We identify some limitations and important future directions of this work, listed below.
Hardware improvement. Common failure cases, such as spherical and ball-shaped objects, are extremely challenging with a parallel gripper, and moving to a multi-fingered hand is an important future direction. Even for humans, retrieving buried objects with only two fingers is a challenging task, putting a realistic upper bound on the performance with parallel grippers. In addition, with a multi-fingered hand, we can explore more complicated action spaces and action primitives, such as scooping.
More complicated action space. Our method restricts the gripper movement inside the granular media to a 2D plane, and future work includes exploring more complicated action spaces (such as 6DOF grasp at an angle) and diverse action primitives (such as scooping, especially with multi-fingered hands). The expanded action space will also require a more efficient training pipeline.
More challenging tasks. We made some assumptions to simplify the task, such as a guaranteed first touch and single object retrieval. In future directions, we want to relieve the assumptions by incorporating initial searching and finding policies to actively locate the target object. We also want to study the retrieval of a particular object among many buried objects. These are more challenging and similar to real-world scenarios, and will likely require a combination of RL training and additional heuristic guidance.
Generalization to other granular types. Our experiments involve one type of granular media, and in our future work, we aim to evaluate our method on other types of granules. We note, however, that our beads have relatively large, non-spherical shapes and display anisotropic behavior. Sands and soil, which have much smaller granules and more isotropic properties, might make it easier to distinguish pushing contact from ubiquitous contact. Other failure modes, such as objects sliding out of a grasp due to a layer of granules stuck between fingers and objects, might also be less frequent. It is also worth noting that smaller granules have less compressibility, potentially generating higher resistance forces for both ubiquitous contact and pushing contact. This necessitates the recalibration of the force threshold to filter ubiquitous contacts.
We also speculate that larger irregular gravel or heterogeneous granular media will be more challenging than the granular beads that we tested on. There will be more jamming for larger or irregular gravels, and the pushing contact will be more difficult to distinguish from ubiquitous contact, meaning a larger gap between open-air policy and granular media policy. In fact, if we keep increasing the size of the granules, such that the granules and the object are of similar size, then the task will become more of a targeted object singulation task. Of course, these hypotheses will need to be rigorously tested in future work.
7 Conclusion
@@
In this work, we propose GEOTACT, a novel method that reliably retrieves objects buried completely under granular media using only touch sensing. This is a challenging task due to the absolute absence of vision and spurious and noisy tactile readings, introduced by the ubiquitous contacts inside granular media. We formulate this task as a model-free reinforcement learning problem and train it end-to-end in simulation with a curriculum strategy. We first train the policy in an open-air environment and then fine-tune it inside a granular environment. Our formulation of the learning task action space leads to emergent pushing behaviors that help reduce uncertainty and also move the object to form a stable grasp. Compared to heuristic-based methods, our learned policy can incorporate simulated sensor noise during training and is thus more robust to spurious and noisy tactile readings. Despite being trained only on 7 objects with simple shapes in simulation, we show that our method can zero-shot transfer to the real hardware and successfully pick up 28 unseen objects, including rigid, deformable, and articulated objects with various complex shapes.
[…]
@@
Acknowledgements We would like to thank Eric Chang and Amr El-Azizi for their support on the DISCO tactile finger. We would like to thank Zhenjia Xu for his help with the granular media simulation. This material is based upon work supported by the National Science Foundation under Award No. CMMI-2037101, and the Office of Naval Research under grant N00014-21-1-4010. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation or the Office of Naval Research.